### v1

# From Intuition to Expertise: Rubric-Based Cognitive Calibration for Human Detection of LLM-Generated Korean Text

Abstract
Distinguishing human-written Korean text from fluent LLM outputs remains difficult even for linguistically trained readers, who can over-trust surface well-formedness. We study whether expert detection can be treated as a learnable skill and improved through structured calibration. We introduce LREAD, a rubric derived from national Korean writing standards and adapted to target micro-level artifacts (e.g., punctuation optionality, spacing behavior, and register shifts). In a three-phase longitudinal blind protocol with Korean linguistics majors, Phase 1 measures intuition-only detection, Phase 2 enforces criterion-level scoring with explicit justifications, and Phase 3 evaluates domain-focused mastery on held-out elementary essays. Across phases, majority-vote accuracy increases from 60% to 100%, accompanied by stronger inter-annotator agreement (Fleiss’ $\kappa$: -0.09 $\rightarrow$ 0.82). Compared to state-of-the-art LLM detectors, calibrated humans rely more on language-specific micro-diagnostics that are not well captured by coarse discourse priors. Our findings suggest that rubric-scaffolded expert judgment can serve as an interpretable complement to automated detectors for non-English settings, and we release the full rubric and a taxonomy of calibrated detection signatures.
From Intuition to Expertise: Rubric-Based Cognitive Calibration for Human Detection of LLM-Generated Korean Text
footnotetext: $\dagger$ Corresponding author.
1 Introduction
The proliferation of large language models (LLMs) has reshaped digital communication and intensified concerns about academic (perkins2023academic; peng-etal-2023-hidding; zhu2024embracing) and creative integrity (yang-etal-2024-survey; wu-etal-2025-survey). Recent models also generate text that is highly fluent and norm-conforming, which can appear authoritative even to trained readers. Although prior work reports systematic differences between human and machine writing, such as increased dependency depth in machine-generated outputs alongside greater semantic variety in human text (zanotto-aroyehun-2025-linguistic), these cues do not reliably translate into stable human judgments. In practice, human annotators often over-rely on surface well-formedness, as polished style and local grammaticality can obscure subtle structural irregularities, resulting in what we characterize as a persistent fluency trap for linguistically trained yet task-uncalibrated readers.
A common expectation is that domain expertise should reduce this vulnerability. Evidence from English settings supports this view: frequent LLM users can outperform commercial systems such as GPTZero [^1] https://gptzero.me/ when identifying synthetic articles (russell-etal-2025-people). At the same time, the majority of detection evidence and tooling remains English-centered, while high-stakes misuse increasingly occurs across diverse languages and writing norms. Non-English settings introduce additional sources of ambiguity, including language-specific morphology, orthography, and punctuation conventions, which can alter both the appearance of fluency and the distribution of artifacts. These factors motivate language-specific, human-centered detection studies that characterize expert reasoning in practice and promote consistent and explainable judgments.
We focus on Korean as a representative case of a morphologically rich language with flexible spacing conventions. Within the Korean linguistic context, prior work identifies specific morphological and punctuational artifacts, including distinctive comma usage patterns, that distinguish LLM-generated Korean text from human-written Korean text (park-etal-2025-katfishnet). Our preliminary findings reveal a distinct failure mode for intuitive detection in this setting. Even Korean linguistics majors, despite formal training in grammar, discourse, and style, achieve only 60% accuracy in a blind attribution task. Their decisions further exhibit low initial inter-annotator agreement, suggesting not only the presence of expert error but also the absence of a shared evaluative criterion. This pattern aligns with the broader observation that humans struggle to draw stable boundaries between human- and machine-authored text, yet can improve through iterative feedback and structured intervention (dugan2023real). We therefore hypothesize that the primary bottleneck is the absence of an explicit and transferable inferential procedure, rather than an inherent deficit in linguistic competence.
We introduce the linguistically-regularized expert AI detection rubric (LREAD), a diagnostic framework derived from national linguistic standards for systematic textual analysis. LREAD functions as a cognitive scaffold that externalizes expert intuition into an explicit, criterion-level decision protocol requiring concrete evidence for each judgment. We integrate LREAD into a three-phase longitudinal blind test that traces the evolution of expert detection from intuition-driven screening to disciplined micro-diagnostics. Across phases, annotators improve detection accuracy from 60% to 100%. This improvement coincides with convergence in judgment, reflected in a marked increase in inter-annotator agreement measured by Fleiss’ $\kappa$. Calibrated experts also converge on a shared lexicon of LLM-generation cues, including cross-lingual interference (e.g., translationese) and model-specific stylistic accents.
2 Intuition and the Detection Gap
2.1 Experimental Setup
We utilize the KatFish corpus (park-etal-2025-katfishnet), the first benchmark dataset specifically designed for detecting LLM-generated Korean text. We select argumentative essays as the primary genre for evaluation because they require a complex balance of logical reasoning and stylistic consistency. Our evaluation set comprises 30 samples selected through stratified random sampling from elementary, middle, and high school levels. Each level contains ten essays, consisting of two human-authored pieces and eight LLM-generated outputs. The LLM-generated portion includes two samples each from GPT-4o, Solar-10.7B, Qwen2-72B, and Llama3.1-70B.
We recruit three undergraduate students majoring in Korean Language and Literature to serve as human annotators. These annotators demonstrate high linguistic sensitivity while having minimal prior exposure to diverse large language models. During the initial blind test phase, we intentionally refrain from providing any formalized rubric or explicit decision criteria, allowing annotators to rely solely on their intuitive linguistic judgment. Each annotator is presented with one essay at a time and is instructed to determine whether the text was authored by a human or generated by an AI system.
For the automated baseline, we employ three state-of-the-art LLMs: GPT-5.2 Thinking, Gemini 3 Flash, and Claude Sonnet 4.5. We provide all participants with the same essay samples and instructions to perform a blind classification task without prior information regarding label distribution or source models.
Evaluator Participant Accuracy (%)
Human Annotator 1 50.00
Annotator 2 46.67
Annotator 3 80.00
Human Majority Vote 60.00
LLMs GPT-5.2 Thinking 96.67
Gemini 3 Flash 96.67
Claude Sonnet 4.5 90.00
LLMs Majority Vote 96.67
Table 1: Detection accuracy for human annotators and LLM detectors in Phase 1.
2.2 Human-LLM Detection Gap
We present the comparative results of the initial blind test in Table 1. Our findings reveal a substantial gap between intuition-driven human attribution and LLM-based reviewing under identical instructions. On this benchmark, LLM detectors achieve comparatively high zero-shot accuracy, whereas human linguistics experts are less reliable under intuition-only attribution.
We identify a 36.67%p gap between the human and LLM majority votes. These results demonstrate that the superficial fluency of modern LLMs successfully bypasses the intuitive linguistic filters of human experts. We find that linguistic expertise does not inherently confer robustness to surface-plausibility bias; experts can conflate technical well-formedness with human authorship and underestimate model-specific artifacts.
2.3 Analysis: Initial Human Perception
We summarize the heuristics that human experts use in the initial blind test to distinguish LLM-generated essays from human writing. These observations reflect the subjective filters that annotators apply before they access ground-truth labels or the formal rubric. We organize these heuristics around structural regularity and discourse plausibility.
Category Evaluation Dimension Points
Content (40) Thesis Clarity 10
Evidence Adequacy and Sufficiency 15
Consideration of Counter-arguments 5
Creativity 10
Organization (10) Global Paragraph-Level Organization 5
Sentence-to-Sentence Cohesion 5
Expression (50) Lexical Usage 10
Sentence Construction 10
Orthographic and Grammatical Conventions 10
Writing Conventions 10
Creativity 10
Total 100
Table 2: Structured overview of the LREAD rubric with point allocation. Detailed scoring criteria for each dimension are provided in the Appendix A.
Annotators treat structural rigidity and mechanical consistency as primary signals of machine generation. They also maintain unusually regular surface patterns, including repetitive sentence openings and overly uniform punctuation, and they exhibit formatting artifacts in mixed-script contexts. Annotators further report that model-written essays adhere to the prompt with excessive literalness, yielding arguments that are normatively safe yet substantively shallow, and thereby constraining perspective-taking. Annotators also interpret conservative lexical choices and low semantic risk-taking as consistent with machine-generated text.
In contrast, annotators associate human authorship with linguistic fallibility and creative deviation. Human essays exhibit minor orthographic errors, localized grammatical inconsistencies, and idiosyncratic metaphors, yielding prose that is less polished yet more individual in voice. Annotators also perceive a more organic rhetorical progression in human writing, including personal anecdotes and emotionally grounded appeals. Overall, human annotators initially equate technical polish with artificiality and treat controlled inconsistency as a cue of human identity.
3 LREAD: A Human-Centered Cognitive Calibration Framework
3.1 Rubric Design from Linguistic Standards
We derive the LREAD rubric from the 2025 standards for argumentative writing proficiency issued by the National Institute of Korean Language. While these national standards provide a robust foundation for general writing assessment, we reconfigure the criteria to heighten sensitivity toward computational artifacts. Importantly, we design the rubric after the completion of the Phase 1 blind evaluation, once the three expert annotators were informed of the ground-truth labels. This post-hoc design choice ensures that the rubric systematically encodes the diagnostic cues that human experts identified during unconstrained judgment. We transform the standard assessment into a specialized detection tool by implementing two fundamental modifications.
Phase Cognitive Mode Intervention & Activity Evaluation Focus
Phase 1 Intuitive
Zero-shot detection task using
unstructured linguistic intuition.
Quantifying baseline susceptibility
to surface-fluency bias.
Phase 2 Analytical
Application of LREAD rubric;
eliciting rubric-scaffolded reasoning.
Measuring accuracy gains and
systematicity in detection.
Phase 3 Mastery
Detection on challenging domains using
internalized LREAD criteria.
Verifying domain-specific mastery
and expert-level consensus.
Table 3: Summary of the three-phase cognitive calibration framework, outlining the experimental design and evaluation targets for each calibration stage.
First, we incorporate a dedicated creativity dimension into the content category. Human writing frequently contains idiosyncratic metaphors, personal narratives, and non-statistical semantic leaps that depart from probabilistic expectations. In contrast, LLMs rely on the statistical recombination of training data, which often results in predictable and generic argumentation. We define creativity in this context as the presence of unique human signatures that violate the typical distribution of LLM outputs.
Second, we refine the scoring granularity to capture minute linguistic variances. Because modern LLMs achieve near-perfect grammatical accuracy, standard evaluation metrics often fail to differentiate synthetic text from human prose. Our granular system requires human experts to examine subtle patterns in lexical choice, morphological consistency, and sentence-level cohesion. By shifting the focus from broad fluency to specific structural regularities, we ensure that evaluators remain attentive to the subtle traces of machine intervention.
Table 2 summarizes the structured allocation of evaluation dimensions and corresponding point weights in the LREAD rubric. The rubric assigns a total of 100 points across content, organization, and expression, with expression accounting for the largest proportion. This asymmetric weighting reflects our empirical observation, derived from the Phase 1 blind study, that contemporary LLMs consistently achieve high scores on surface-level coherence and structural adequacy, while fine-grained expressive variation continues to serve as a robust indicator of human authorship.
Notably, the rubric evaluates creativity at both the content and expression levels. This dual placement enables the framework to capture human originality manifested through unconventional argumentative development as well as stylistic deviations from distributionally typical language patterns. By operationalizing creativity across multiple linguistic strata, the LREAD rubric equips annotators to identify subtle yet systematic traces of machine generation that remain undetected by automated detectors.
3.2 Three-Phase Cognitive Calibration
We propose a three-phase longitudinal calibration framework designed to trace and induce the evolution of human detection expertise. Rather than treating human judgment as static, this framework models detection as a learnable cognitive skill that emerges through iterative feedback and structured analytical scaffolding. The primary objective of the proposed framework is to systematically transform linguistically trained yet task-uncalibrated annotators into calibrated expert detectors, enabling them to reliably identify latent computational signatures that extend beyond surface-level linguistic fluency. The overall trajectory of this framework, detailing the transition of cognitive modes and evaluation targets at each stage, is summarized in Table 3.
Phase 1 establishes an intuitive baseline through a zero-shot blind detection task, as described in Section 2. During this phase, annotators rely solely on unstructured linguistic intuition, without access to a formal rubric or explicit decision criteria. We use this phase to quantify experts’ initial susceptibility to surface-fluency bias, whereby stylistic polish in LLM-generated text obscures underlying structural artifacts. Following Phase 1, we disclose the ground-truth labels and the identities of the four source models. This disclosure serves as an initial calibration signal, enabling annotators to retrospectively align their intuitive judgments with model-specific output patterns.
Phase 2 introduces structured analytical detection. Building on the insights gained from Phase 1, annotators collaboratively design the LREAD rubric by integrating national writing standards with empirically observed failure cases. We then conduct a second blind evaluation using a new set of 30 essay samples, stratified evenly across elementary, middle, and high school levels. We construct these samples to be disjoint from those used in Phase 1 and deliberately withhold the label distribution to mitigate memorization effects and distributional inference. During this phase, annotators justify each classification decision by providing explicit, multi-dimensional scores grounded in the LREAD rubric. This design elicits rubric-scaffolded reasoning, prompting evaluators to align perceived linguistic irregularities with formal analytic categories rather than relying on holistic impressions.
Phase 3 represents domain-specific mastery. After a second disclosure of ground-truth labels, annotators perform a final blind detection task focused exclusively on elementary-level essays. We deliberately restrict this phase to the elementary domain, as it constitutes the most challenging setting for generation fidelity. We restrict this phase to the elementary domain, which poses the most stringent test of generation fidelity. Our evaluation set contains 10 essays: 2 human-written and 8 LLM-generated (2 each from GPT-4o, Solar, Qwen2, and Llama3.1), sampled using the same stratified strategy as in Phase 1. valentini-etal-2023-automatic show that state-of-the-art LLMs struggle to control lexical simplicity and readability when generating children’s stories, often exceeding age-appropriate vocabulary constraints. Complementing this finding, hassan2025evaluating demonstrate that expert evaluators systematically perceive LLM-generated child-like text as linguistically more advanced than the target age group. More broadly, wang-etal-2025-catch find that LLMs have difficulty imitating implicitly defined writing styles, tending to default to structurally over-regularized and stylistically polished outputs rather than capturing subtle, informal authorial signals. In the final phase, annotators apply the internalized rubric to detect persona-mismatch artifacts. We conclude by eliciting from calibrated experts explicit accounts of each source model’s linguistic fingerprints, alongside defining features of authentic human writing.
3.3 Experimental Procedure
We enforce strict information control throughout the three-phase calibration process to preserve experimental integrity. At no point do we disclose the proportion of human-authored versus LLM-generated essays, thereby preventing statistical inference or prior-driven guessing. Each phase employs entirely distinct text samples to eliminate memory-based recognition and ensure that performance gains reflect genuine cognitive calibration rather than familiarity effects.
The transition from Phase 1 to Phase 3 is designed as a cognitive progression: annotators move from intuitive judgments to rubric-guided analysis, and ultimately to domain-specific expertise. This design enables us to test how rubric scaffolding and iterative feedback convert general linguistic competence into specialized detection skill.
By integrating longitudinal exposure, controlled feedback, and analytical justification, our framework operationalizes cognitive calibration as a measurable and reproducible process. This design supports fine-grained analysis of how human experts move beyond superficial fluency to attend to stable, interpretable signals of machine generation.
4 Results of Human Calibration
4.1 Accuracy Gains Across Calibration Phases
Figure 1 tracks how detection performance evolves under the three-phase calibration protocol. In Phase 1, annotators rely on intuition alone and the human majority vote reaches 60.00% accuracy, which trails the LLM-detector majority vote (96.67%) by 36.67%p. This gap indicates that fluent LLM outputs reliably exploit surface-level plausibility cues that human annotators typically trust when they lack structured decision criteria. Phase 2 introduces rubric-guided judgment and shifts detection from impression-based screening to explicitly justified evaluation. The result shows that this intervention substantially improves accuracy for each annotator and for the majority vote (60% $\rightarrow$ 90%). Appendix B further characterizes Phase 2 with generator-wise score distributions and annotator-level descriptive statistics of total LREAD scores. Phase 3 completes the calibration loop and yields perfect majority-vote performance.
Figure 1: Detection accuracy across the three phases for each human annotator and their majority vote.
4.2 Error Dynamics and Calibration
Figure 2 shows how calibration reshapes majority-vote errors. In Phase 1, the majority vote yields no false positives (Human mislabeled as AI) but many false negatives (AI mislabeled as Human): the confusion matrix shows 6 true-human essays correctly classified as Human, 0 true-human essays mislabeled as AI, 12 AI essays mislabeled as Human, and 12 AI essays correctly labeled as AI. This pattern suggests that surface-level fluency systematically cues human authorship, causing annotators to mislabel many highly fluent LLM-generated essays as human-written.
(a) Phase 1
(b) Phase 2
Figure 2: Confusion matrices for the human-annotator majority vote in Phase 1 (intuition-only) and Phase 2 (rubric-guided). Each cell reports the number of essays in the corresponding true-label and predicted-label pair under the binary Human versus AI decision.
In Phase 2, rubric guidance sharply reduces false negatives and reallocates errors toward a smaller number of false positives. The confusion matrix records 5 true-human essays correctly classified as Human and 1 mislabeled as AI, while it labels 22 AI essays correctly as AI and misclassifies only 2 as Human. This shift indicates that the rubric successfully directs attention away from superficial polish and toward more stable diagnostic cues, improving recall on AI while introducing only limited over-correction on human essays.
Figure 3 complements the confusion-based error analysis by visualizing Phase 2 rubric score distributions across generators. By aggregating criteria into Content, Organization, and Expression, the plot frames LREAD as a structured, multidimensional competency profile and reveals systematic source-level variation across dimensions and models.
Figure 3: Phase 2 rubric profile by generator. The radar chart reports the mean macro scores (Content, Organization, Expression) aggregated from the rubric dimensions and normalized by each category maximum (40/10/50).
4.3 Inter-annotator Agreement Analysis
Table 4 reports Fleiss’ $\kappa$ across phases using the inclusive definition that maps any non-binary response to a third category. We use this formulation because the protocol allows an explicit uncertainty response. This choice aligns with our goal of measuring calibration as a cognitive process: uncertainty is an observable signal in early phases, and the inclusive mapping retains it rather than discarding the corresponding items.
The agreement trajectory mirrors the accuracy gains. Phase 1 yields negative agreement ($\kappa=-0.09$), which indicates systematic inconsistency under intuition-only detection. Phase 2 increases agreement to $\kappa=0.24$, reflecting that rubric-guided criteria align annotators toward shared decision boundaries. Phase 3 reaches strong agreement ($\kappa=0.82$) on the calibrated test set, consistent with a converged expert representation that supports both high accuracy and stable consensus.
Phase N Fleiss’ $\kappa$
Phase 1 (Intuition-only) 30 -0.09
Phase 2 (Rubric-guided) 30 0.24
Phase 3 (Calibrated) 10 0.82
Table 4: Inter-annotator agreement (Fleiss’ $\kappa$) across phases (N = essays per phase). We report the inclusive variant that maps any non-binary response to a third category, preserving a consistent item set while treating abstention or uncertainty as an explicit label.
5 Machine Perception vs. Human Insight
We compare the calibrated human signatures with signatures produced by three LLM-based reviewers after Phase 1. This contrast isolates which cues are salient to strong language models under zero-shot inspection and which cues become diagnostic through human calibration. We provide the consolidated, model-wise human signature inventory in Appendix C and focus here on the comparative analysis with LLM detectors.
5.1 Human–LLM Comparison Protocol
The comparison contrasts two qualitatively different sources of evidence. The human signatures were elicited after Phase 3, after annotators received gold labels for all phases and completed rubric-based calibration. The LLM signatures were produced after Phase 1, before rubric construction, reflecting zero-shot inspection by strong language models. This setup helps disentangle broadly accessible cues from calibration-specific micro-diagnostics that emerge through iterative feedback.
Cue Family Human Annotators LLM Detectors Relation
Macro-structure Model-specific templates (e.g., enumeration and conclusion openings); template mixing becomes diagnostic for some models. Emphasis on canonical 3-part structure and balanced concession patterns; “too complete” structure as an AI cue. Overlap
Surface mechanics Optional-but-recurrent comma placement around connectives and clause endings; spacing irregularities; within-essay inconsistency as a cue. Mechanics used mainly as a coarse cue: near-zero typos implies AI; disfluency clusters imply human. Divergence
Lexicon and register Sino-Korean density and nominalization as discriminators; two types of translationese identified—Sino-Korean calques (Qwen2) versus English-style modifier-noun constructions (GPT-4o). High-level register labels (report-like tone, mature vocabulary) without pinpointing micro-realization patterns. Partial
Persona and affect Writer persona becomes diagnostic; collective responsibility and forward-looking closures recur as AI cues. Emotion and lived experience as human cues; safe neutrality and socially agreeable resolutions as AI cues. Overlap
Evidence style Concrete elaboration and counterargument handling differentiate models; generic remedies appear suspicious. Personal anecdotes and messy constraints imply human; generic institutional prescriptions imply AI. Overlap
Table 5: Comparison of cues emphasized by calibrated human annotators and LLM detectors. “Relation” indicates whether the cue family primarily overlaps or diverges in operational granularity.
5.2 Translationese: A Human Detection Signal
As summarized in Table 5, the contrast between calibrated human annotators and LLM-based reviewers reveals both substantive overlap and systematic divergence across cue families. At a high level, both groups attend to macro-structural regularities and generic resolution patterns, suggesting that certain discourse-level signals of machine authorship are broadly accessible even under zero-shot inspection. However, the table also highlights a consistent asymmetry in operational granularity: calibrated humans attend to stylistically optional but model-specific punctuation habits, such as the excessive use of commas after sentence-initial connectives or clause-linking endings, which LLM detectors frequently overlook as surface fluency. This divergence motivates the subsequent analysis, which distinguishes shared cues that emerge without calibration from translationese-like and punctuation-level phenomena that become discriminative only through iterative human feedback.
5.3 Shared Cues
Humans and LLM evaluators broadly converged on macro-regularity as a primary signal of machine authorship. Both groups identified a highly standardized “General-to-Specific” logic flow and a mechanical 3-part structure, particularly in GPT-4o, which consistently utilizes ordinal markers (e.g., “First”, “Second”) to partition arguments. Both also treated repetition—either at the level of connectives (Solar) or at the level of sentence-final realizations—as a strong cue, and both observed that LLM essays often terminate in broadly agreeable prescriptions (e.g., education or institutional reform) rather than topic-specific, lived-experience constraints.
5.4 Divergent Cues
The primary divergence concerns granularity. Following calibration, human annotators relied on Korean-specific micro-patterns that resist reduction to generic fluency judgments, including optional comma placement after sentence-initial connectives, punctuation following clause-linking endings, spacing at multiword boundaries, and register diagnostics associated with Sino-Korean nominalization. In contrast, LLM detectors emphasized global properties such as overall coherence, balanced dialectical framing, and the safety of examples and solutions. Although these properties often reflect shared underlying regularities, they operate at a coarser granularity and do not explicitly encode language-specific mechanisms.
Human annotators consistently observed a translated quality in model outputs, with Qwen2 displaying Sino-Korean nominalization and calque-like phrasing, and GPT-4o exhibiting English-style translationese characterized by modifier-heavy noun phrases and frequent pronominal reference. Both patterns were grammatically well-formed yet rhetorically non-native. Prior work has shown that translationese carries systematic source-language fingerprints and stylistic shifts that can make translated text distributionally distinct from natively authored text (doshi-etal-2024-pretraining). Evaluation pipelines that rely on translated instructions can further confound naturalness, complicating fair comparisons between English-conditioned and target-language behaviors (enomoto-etal-2025-fair).
Recent studies connect these artifacts to multilingual LLM training and alignment. li-etal-2025-lost show that supervised training can shape literal, translation-like realizations, and guo-etal-2025-large report an “English accent” effect in multilingual generation that degrades perceived target-language naturalness. We therefore treat translationese-like interference as a discriminative cue family that calibrated humans can exploit, and we provide an expanded discussion with concrete measurement candidates in Appendix D.
6 Related Work
6.1 Post-hoc Detection & LLM Watermarking
Prior work on attributing LLM-generated text has largely followed two directions: post-hoc detection (mitchell2023detectgpt; su-etal-2023-detectllm; bao2024fastdetectgpt; zhu-etal-2023-beat; yang2024dnagpt) and LLM watermarking (kirchenbauer2023watermark; dathathri2024scalable; lu-etal-2024-entropy; chen-etal-2024-watme; hou-etal-2024-semstamp; guo2024context; wang-etal-2025-morphmark; park2026watermod).
Post-hoc detection methods typically operate without altering the source model, making them suitable for retrospective auditing. DetectGPT (mitchell2023detectgpt) exploits probability curvature in the likelihood landscape, with subsequent work improving efficiency via log-rank–based estimators (su-etal-2023-detectllm) and conditional probability curvature (bao2024fastdetectgpt). Other approaches leverage LLM querying through revision-based similarity analysis (zhu-etal-2023-beat) or analyze divergent continuations and n-gram distributions (yang2024dnagpt). Despite strong empirical performance, these methods remain sensitive to domain shift and continued advances in generation quality.
LLM watermarking approaches address these limitations by embedding attribution signals directly into the decoding process. kirchenbauer2023watermark introduce a foundational logit-bias watermarking framework, which subsequent work extends toward scalable deployment (dathathri2024scalable) and a range of design variants, including entropy-aware (lu-etal-2024-entropy), lexical (chen-etal-2024-watme), semantic (hou-etal-2024-semstamp), and contextual (guo2024context) schemes. MorphMark (wang-etal-2025-morphmark) advances this line by adaptively modulating watermark strength using the cumulative probability mass of green-list tokens to address the effectiveness–quality trade-off. WaterMod (park2026watermod) refines the paradigm through probability-ranked modular partitioning. Collectively, these approaches establish strong baselines while treating attribution as automated inference, with limited emphasis on human judgment.
6.2 Linguistic Profiling of AI-Generated Text
Recent work has shifted from purely algorithmic detection toward systematic linguistic characterization of LLM-generated text. Rather than framing detection as a black-box task, this line of research examines how model outputs differ from human writing across multiple levels of linguistic structure.
zanotto-aroyehun-2025-linguistic find that LLM-generated text often shows deeper and longer dependency structures than human writing, while human texts exhibit more diverse semantic content and greater overall variation. They further note that newer models become more human-like on syntactic depth and length metrics, shrinking earlier gaps. Within the Korean context, park-etal-2025-katfishnet identify language-specific artifacts in spacing, morphology, and punctuation that differentiate LLM-generated Korean text from human writing. These observations highlight the importance of attending to fine-grained linguistic realizations beyond surface-level fluency.
6.3 Human Expertise in AI Detection
Recent studies increasingly emphasize the role of expertise, experience, and structured feedback in shaping human detection performance. Rather than treating human judgment as inherently unreliable, this work shows that detection ability varies substantially with exposure and analytic framing.
russell-etal-2025-people demonstrate that frequent LLM users outperform many automatic detectors when identifying synthetic English articles, while providing interpretable rationales grounded in linguistic cues. dugan2023real demonstrate that boundary detection remains challenging and exhibits substantial inter-annotator variability; nevertheless, targeted interventions are shown to produce measurable improvements over time.
7 Conclusion
We demonstrate that human attribution of LLM-generated text is a learnable and calibratable competency, not a fixed bottleneck of expert judgment. In a three-phase longitudinal study, linguistically trained annotators transition from surface-fluency heuristics to LREAD-scaffolded analysis, achieving substantial accuracy gains with markedly higher inter-annotator agreement and improved sensitivity to language-specific signals that automated detectors still struggle to capture reliably.
Limitations
Annotator Scale and Homogeneity.
This study relies on three annotators with homogeneous training in Korean linguistics, which limits claims about population-level detection accuracy and generalization to non-expert or cross-disciplinary settings. We acknowledge that with only three raters and 10 items in Phase 3, our study is statistically underpowered to detect small effects or to rule out ceiling artifacts; the 100% majority-vote accuracy in Phase 3 should therefore be interpreted as an upper-bound demonstration rather than a robust population estimate.
At the same time, this constrained design reflects a deliberate methodological choice. By recruiting annotators with matched linguistic backgrounds and minimal prior LLM exposure, we isolate the causal contribution of cognitive calibration from confounds introduced by heterogeneous expertise or divergent annotation strategies. The resulting analysis prioritizes internal validity and theoretical clarity, establishing calibration as a learnable, structured property of human judgment rather than a demographic or individual-difference effect. Future work should expand this protocol to larger and more diverse annotator pools—including non-experts and cross-linguistic populations—to quantify how calibration dynamics interact with baseline expertise, prior LLM familiarity, and task incentives.
Genre and Domain Scope.
The empirical evaluation focuses exclusively on argumentative essays at the primary and secondary school levels, a genre that foregrounds discourse organization and reasoning coherence. This controlled setting enables precise measurement of how rubric scaffolding reshapes judgments beyond surface fluency, but it leaves open whether the same calibration trajectory holds for genres with different stylistic conventions—conversational dialogue, narrative fiction, or technical documentation. Extending LREAD to these domains will require principled rubric adaptation and systematic revalidation.
Post-hoc Rubric Design.
The LREAD rubric was constructed after Phase 1, once annotators had access to ground-truth labels. While this post-hoc design ensures that the rubric encodes empirically grounded failure modes, it raises potential concerns about overfitting to the specific samples used in rubric development. We mitigate this risk by using entirely disjoint essay sets across phases and withholding label distributions throughout; nevertheless, independent validation on held-out corpora remains necessary to confirm generalizability.
Model and Temporal Generalization.
Our linguistic diagnostics are derived under fixed model versions (GPT-4o, Solar-10.7B, Qwen2-72B, Llama3.1-70B) and a single generation condition. As LLM capabilities evolve, some cues that currently discriminate machine-generated text—such as particular punctuation regularities or translationese patterns—may diminish or shift. Longitudinal and cross-model studies will be essential for distinguishing stable, transferable cues from transient artifacts tied to specific model generations, and for evaluating the durability of calibration effects as the detection landscape changes.
Ablation of Calibration Components.
Our three-phase protocol bundles multiple interventions: gold-label feedback, collaborative rubric design, and explicit scoring requirements. The current study does not isolate the contribution of each component; future work should include ablation conditions (e.g., feedback-only without rubric, rubric without gold-label disclosure) to clarify which elements drive the observed accuracy and agreement gains.
Appendix A LREAD Rubric Specifications
We include the complete Korean LREAD rubric in Figure 4 and an English translation in Figure 5. The rubric specifies anchored scoring criteria across three dimensions: Content (40), Organization (10), and Expression (50), with discrete score levels for each sub-criterion. Percentage-based anchors follow unit-normalized counting rules: lexical-level issues are measured as a proportion of eojeols (Korean spacing units), sentence-level issues are measured as a proportion of sentences, and orthographic-norm violations are measured as a proportion of words. We apply conservative deduplication when the same error pattern recurs, counting repeated instances of an identical grammatical phenomenon (e.g., particle misuse) as a single case, and treating repeated spelling errors with the same erroneous form as a single case. When a judgment requires contextual interpretation, annotators document the decision with a representative example excerpt to support auditability and consistent application.
Figure 4: The full LREAD rubric with anchored scoring criteria (Korean original).
Figure 5: The full LREAD rubric with anchored scoring criteria (English translation).
Appendix B Phase 2 Rubric Score Distributions and Annotator-Level Descriptives
This appendix complements the main Phase 2 results by (i) visualizing generator-wise distributions of total rubric scores and (ii) reporting annotator-level descriptive statistics that contextualize grade-level and generator-level variability under the same scoring protocol.
B.1 Total LREAD Score Distributions
[…]
Figure 6 summarizes the distribution of total LREAD scores in Phase 2 by generator (Human, GPT-4o, Solar, Qwen2, and Llama3.1). The figure pools rater–essay evaluations across three annotators, which characterizes both central tendency and dispersion for each source. Each generator contributes six essays (two per grade level), yielding 18 pooled evaluations per generator (6 essays $\times$ 3 annotators).
While the main paper focuses on binary detection outcomes, total LREAD scores provide a graded view of rubric-guided judgments. Higher totals indicate stronger satisfaction of the rubric’s content, organization, and expression criteria, whereas lower totals reflect accumulated deficiencies revealed by the analytic protocol. Generator-wise distributions therefore provide complementary evidence about rubric-guided auditing beyond surface-level fluency. In particular, tighter dispersion is consistent with more uniform rubric outcomes across prompts and grade strata, whereas broader dispersion is consistent with greater variability in rubric-aligned writing quality.
Evaluator Elementary Middle High
Annotator 1 63.60 $\pm$ 19.59 72.90 $\pm$ 12.89 75.40 $\pm$ 8.54
Annotator 2 60.10 $\pm$ 13.77 72.00 $\pm$ 10.92 69.40 $\pm$ 7.72
Annotator 3 63.20 $\pm$ 13.18 68.60 $\pm$ 13.83 67.90 $\pm$ 5.78
All (pooled) 62.30 $\pm$ 15.31 71.17 $\pm$ 12.31 70.90 $\pm$ 7.90
Table 6: Phase 2 total LREAD scores by grade level (mean $\pm$ SD). Total scores are summarized for each annotator and for pooled evaluator–essay evaluations. Each grade level includes 10 essays, yielding 30 pooled evaluations per level across three annotators.
Evaluator Human GPT-4o Solar Qwen2 Llama3.1
Annotator 1 50.67 $\pm$ 19.61 76.50 $\pm$ 4.09 74.00 $\pm$ 4.05 82.00 $\pm$ 3.90 70.00 $\pm$ 12.41
Annotator 2 55.33 $\pm$ 15.97 73.67 $\pm$ 4.27 64.67 $\pm$ 9.35 72.50 $\pm$ 10.56 69.67 $\pm$ 8.94
Annotator 3 54.33 $\pm$ 6.28 72.17 $\pm$ 12.45 64.33 $\pm$ 12.08 73.33 $\pm$ 8.33 68.67 $\pm$ 7.63
All (pooled) 53.44 $\pm$ 14.28 74.11 $\pm$ 7.70 67.67 $\pm$ 9.73 75.94 $\pm$ 8.79 69.44 $\pm$ 9.29
Table 7: Phase 2 total LREAD scores by generator (mean $\pm$ SD). Scores are summarized for each annotator and for pooled evaluator–essay evaluations. Each generator contributes 6 essays (2 per grade level), yielding 18 pooled evaluations per generator across three annotators.
B.2 Annotator-Level Descriptive Statistics
Tables 6 and 7 report mean $\pm$ standard deviation (SD) of total LREAD scores in Phase 2. These summaries support two complementary interpretations. First, they report score magnitude and variability for each annotator, which contextualizes the pooled distributions in Figure 6. Second, they provide grade-level and generator-level views of total-score variation under a unified rubric.
Table 6 aggregates scores by grade level (elementary, middle, high) for each annotator and for pooled evaluations. Each grade level contains 10 essays, yielding 30 pooled evaluations per level across three annotators (10 essays $\times$ 3 annotators). These summaries contextualize rubric-guided scoring across developmental writing regimes, which differ in rhetorical maturity and normative writing proficiency.
Table 7 reports statistics by generator, enabling direct comparison between human-authored essays and each LLM source under a unified analytic protocol. Each generator contributes six essays (two per grade level), yielding 18 evaluations across three annotators (6 essays $\times$ 3 annotators). Generator-wise means summarize central tendency, while standard deviations quantify dispersion in rubric scores across prompts and grade levels.
Figure 6 provides a distributional view of total rubric outcomes, whereas Tables 6 and 7 provide complementary descriptive summaries that separate grade-level and generator-level variation. Taken together, these results indicate that rubric-guided auditing produces comparable scoring patterns across sources within an explicit analytic framework, while clarifying whether observed variation reflects writing level rather than generator identity.
B.3 Creativity Diagnostics
The LREAD rubric includes two creativity-related criteria, one under Content and one under Expression. Because creativity scores are sparse (many evaluations receive zero), we report mean $\pm$ SD and the proportion of non-zero scores (%>>0), which reflects creativity cues under the rubric.
Generator Content %>>0 Expression %>>0
Human 4.44 $\pm$ 5.11 44.40 1.39 $\pm$ 2.30 27.80
GPT-4o 0.56 $\pm$ 2.36 5.60 0.56 $\pm$ 1.62 11.10
Solar 0.56 $\pm$ 2.36 5.60 0.28 $\pm$ 1.18 5.60
Qwen2 2.78 $\pm$ 4.61 27.80 1.39 $\pm$ 2.87 22.20
Llama3.1 1.67 $\pm$ 3.83 16.70 0.33 $\pm$ 1.19 11.10
Table 8: Phase 2 creativity-related rubric scores by generator (mean $\pm$ SD and %>>0). The table reports pooled evaluator–essay evaluations across three annotators. Content and Expression correspond to the creativity criteria under the LREAD rubric. %>>0 denotes the percentage of evaluations with a non-zero creativity score.
Table 8 indicates higher incidence of non-zero creativity scores for human-authored essays than for most LLM sources, particularly on the content-side criterion. Some LLMs (e.g., GPT-4o and Solar) rarely trigger these creativity cues under the rubric, whereas other models exhibit a higher but still lower-than-human incidence. These patterns align with rubric-guided evaluation that rewards localized deviation and idea-level novelty beyond uniformly fluent realization.
Appendix C Taxonomy of Calibrated Human Detection Signatures
We synthesize model-wise fingerprints articulated by human annotators after completing Phase 3 and receiving gold labels for all phases. We treat these reflections as calibrated qualitative evidence of which cues became operational through iterative feedback and rubric-guided reasoning, rather than as claims about intrinsic or universal model traits. The goal is to document which linguistic regularities emerged as diagnostic for humans within our controlled essay evaluation setting.
Following the analytic framework introduced in Section 5, we organize the signatures along two axes: form-level cues—including macro-structure, connective usage, punctuation, and sentence rhythm—and content-level cues—including stance strength, evidential style, and preferred resolution patterns. This bifurcation instantiates the distinction between surface-level mechanisms and discourse-level priors that grounds the human–LLM comparison in Table 5.
C.1 Signature Elicitation Protocol
Annotator notes were coded into a shared cue taxonomy through iterative consolidation. We prioritized patterns that (i) were independently reported by multiple annotators or (ii) sharply differentiates models under identical prompting and scoring regimes. Crucially, we report these signatures as after-calibration descriptors: they capture the cues that become salient once annotators are provided with gold labels and explicit rubrics, rather than impressions formed during pre-calibration or zero-shot inspection.
Model Form-Level Cues (Structure, Rhythm, Mechanics) Content-Level Cues (Stance, Evidence, Resolution)
GPT-4o Strongly standardized essay scaffold: hook $\rightarrow$ pivot $\rightarrow$ thesis in the introduction; frequent ordinal enumeration in the body; conclusions often begin with conventional discourse markers. Annotators consistently identified optional-but-recurrent comma insertion around sentence-initial connectives and following clause-linking endings. Low-risk and socially agreeable argumentation with restrained stance strength. Essays frequently close with forward-looking prescriptions and projecting generalized future outcomes, often invoking inclusive subjects (e.g., “we”) to suggest directions for society or readers. Unlike Qwen2’s emphasis on institutional duty, GPT-4o frames collective action as aspirational rather than obligatory.
Solar Connective-heavy realization with a narrow connective repertoire and repeated tokens within a single essay. Sentence rhythm tends to be short and uniform, with frequent comma insertion in positions that are grammatically permissible yet stylistically optional. Despite the frequent use of explicit connective devices, local coherence remains uneven and episodic. High surface redundancy with comparatively thin elaboration. Argument flow sometimes oscillates between claim and rebuttal moves, and explicit self-markers of opinion (e.g., “I think”, “my view”) appear performative rather than evidential.
Qwen2 Relatively natural punctuation rhythm; frequent use of formal connectives to construct explicit logical chains; frequent nominalizations and dense Sino-Korean lexicon; occasional spacing errors at multiword boundaries; recurrent quotation-like emphasis around abstract nouns. Annotators consistently identified the presence of lexical calques and rhetorical non-nativeness, localized primarily within syntactically complex modifier-noun constructions. Recurring educational or institutional framing, with conclusions emphasizing collective duty and responsibility. Figurative language appears comparatively often, while the overall register is frequently judged overly mature for the nominal student level.
Llama3.1 Highest stylistic variance and frequent template mixing, weakening a single dominant structural cue. Outputs are comparatively shorter, with variable comma consistency even within the same essay. Hypothetical and counterfactual framing (“if”-clauses) appears frequently, alongside localized grammatical awkwardness. More concrete elaboration and careful counterargument handling than other models, including detailed examples. However, lexical repetition can recur without paraphrase across distant spans. A more vivid writer persona sometimes emerges through rhetorical questions or conversational devices.
Table 9: After-calibration model signatures reported by human annotators following Phase 3.
C.2 Model-Specific Detection Signatures
Table 9 presents the resulting LLM-wise signature matrix for the four source LLMs (GPT-4o, Solar, Qwen2, and Llama3.1). Each row summarizes the dominant form-level and content-level cues that annotators repeatedly associated with a given LLM after calibration. The table provides a compact inventory of operationalized human diagnostics, reflecting both what annotators noticed and which cues they learned to prioritize when distinguishing LLM sources across phases and prompts.
C.3 Shared Detection Signatures
Beyond model-specific fingerprints, Table 9 also reveals several cross-model regularities that annotators consistently treated as signals of machine authorship. First, LLM outputs exhibit unusually stable surface correctness relative to student writing, including fewer orthographic errors and more controlled sentence lengths. Second, outputs tend to flatten grade-level variation: essays nominally generated at elementary, middle, and high-school levels converge toward a similar band in polish and abstraction. Third, many essays converge on generic, institution-facing remedies (e.g., education, awareness campaigns, or system improvements) rather than grounded, personally constrained resolutions.
Appendix D Translationese Diagnostics in Calibrated Human Detection
Expanding on the analysis in Section 5.2, we operationalize translationese as a formal framework to account for the systematic stylistic shifts documented in Table 9, which are most pronounced in Qwen2 and notably present in GPT-4o. We further propose a set of measurement candidates that formalize this qualitative intuition as an operationalized family of cues.
D.1 Motivation and Prior Evidence
We use translationese to denote systematic source-language fingerprints and stylistic shifts that distinguish translated text from natively authored text. Such patterns can be learned and propagated during model pretraining when translationese is present in the training data (doshi-etal-2024-pretraining). Multilingual evaluation pipelines that rely on translated instructions may further confound judgments of naturalness and bias comparisons between English-conditioned and target-language behaviors (enomoto-etal-2025-fair). In multilingual LLMs, supervised training objectives can incentivize overly literal realizations that preserve semantic adequacy while degrading native-like naturalness (li-etal-2025-lost). Relatedly, prior work reports an “English accent” effect in multilingual generation, which reduces perceived target-language naturalness (guo-etal-2025-large).
D.2 Translationese Diagnostics
Grounded in the calibrated human signature inventory (Table 9), we outline a set of translationese-sensitive diagnostics that can be instantiated as measurable features:
  • Nominalization Rate: the proportion of nominalized predicates and abstract nominal constructions, capturing “noun-heavy” discourse packaging characteristic of literal transfer.
  • Sino-Korean Lexical Density: the share of Sino-Korean lexemes and high-register vocabulary, which often co-varies with translated or formalized style.
  • Discourse Deixis and Pronominal Reference: frequencies of demonstratives and pronominal forms (e.g., this/that-type anaphora) that are associated with English-style translated discourse structuring, particularly salient in GPT-4o outputs.
  • Connective–Comma Coupling: co-occurrence patterns between sentence-initial discourse connectives and optional comma insertion, reflecting punctuation regularities that annotators flagged as model-specific.
  • Sentence-Level Naturalness Scoring: complementary global measures of perceived naturalness, following prior work on multilingual naturalness evaluation (guo-etal-2025-large).
These diagnostics are designed to shift translationese from an impressionistic descriptor to a reproducible family of cues. They make explicit how calibrated human detection can leverage language-specific micro-realization patterns not captured by coarse fluency or document-level coherence.