### v2

# From Intuition to Calibrated Judgment: A Rubric-Based Expert-Panel Study of Human Detection of LLM-Generated Korean Text

Abstract
Distinguishing human-written Korean text from fluent LLM outputs remains difficult even for trained readers, who can over-trust surface well-formedness. We present LREAD, a Korean-specific instantiation of a rubric-based expert-calibration framework for human attribution of LLM-generated text. In a three-phase blind longitudinal study with three linguistically trained annotators, Phase 1 measures intuition-only attribution, Phase 2 introduces criterion-anchored scoring with explicit justifications, and Phase 3 evaluates a limited held-out elementary-persona subset. Majority-vote accuracy improves from 0.60 in Phase 1 to 0.90 in Phase 2, and reaches 10/10 on the limited Phase 3 subset (95% CI [0.692, 1.000]); agreement also increases from Fleiss’ $\kappa=-0.09$ to $0.82$. Error analysis suggests that calibration primarily reduces false negatives on AI essays rather than inducing generalized over-detection. We position LREAD as pilot evidence for within-panel calibration in a Korean argumentative-essay setting. These findings suggest that rubric-scaffolded human judgment can complement automated detectors by making attribution reasoning explicit, auditable, and adaptable.
From Intuition to Calibrated Judgment: A Rubric-Based Expert-Panel Study of Human Detection of LLM-Generated Korean Text
footnotetext: $\dagger$ Corresponding author.
1 Introduction
The proliferation of large language models (LLMs) has reshaped digital communication and intensified concerns about academic (Perkins, 2023; Peng et al., 2023; Zhu et al., 2024) and creative integrity (Yang et al., 2024b; Wu et al., 2025). Recent models also generate text that is highly fluent and norm-conforming, which can appear authoritative even to trained readers. Although prior work reports systematic differences between human and machine writing, such as increased dependency depth in machine-generated outputs alongside greater semantic variety in human text (Zanotto and Aroyehun, 2025), these cues do not reliably translate into stable human judgments. In practice, human annotators often over-rely on surface well-formedness, as polished style and local grammaticality can obscure subtle structural irregularities, resulting in what we characterize as a persistent fluency trap for linguistically trained yet task-uncalibrated readers.
A common expectation is that domain expertise should reduce this vulnerability. Evidence from English settings supports this view: frequent LLM users can outperform commercial systems such as GPTZero [^1] https://gptzero.me/ when identifying synthetic articles (Russell et al., 2025). At the same time, the majority of detection evidence and tooling remains English-centered, while high-stakes misuse increasingly occurs across diverse languages and writing norms. Non-English settings introduce additional sources of ambiguity, including language-specific morphology, orthography, and punctuation conventions, which can alter both the appearance of fluency and the distribution of artifacts (Park et al., 2025).
These observations motivate a different question: Is the primary bottleneck in human attribution a lack of linguistic knowledge, or the absence of an explicit diagnostic procedure? We investigate this question in Korean, a setting where spacing conventions, register variation, and punctuation optionality weaken the reliability of coarse fluency-based heuristics. In this sense, Korean constitutes more than a non-English extension; rather, it serves as a valuable test case for human-centered attribution.
To this end, we introduce LREAD, a Korean-specific rubric and calibration protocol for structured human attribution of LLM-generated text. Rather than treating human judgment as fixed, we study whether attribution accuracy, agreement, and auditability improve when linguistically trained annotators move from intuition-only screening to criterion-anchored analysis.
Our contributions are fourfold:
  • We present a three-phase expert-calibration protocol for studying human attribution as a learnable skill under strict information control.
  • We introduce LREAD as a Korean-specific instantiation grounded in national writing standards and adapted for attribution-oriented micro-diagnostics.
  • We provide pilot evidence, in a controlled argumentative-essay setting, that rubric scaffolding improves within-panel accuracy and agreement.
  • We release the full rubric, a calibrated cue taxonomy, and transparent reporting templates to support future work on auditable human-centered evaluation.
2 Related Work
Our work sits at the intersection of three lines of research: (i) automated LLM-generated text detection; (ii) linguistic profiling of LLM outputs; and (iii) human expertise in LLM-generated text detection.
2.1 Post-hoc Detection & LLM Watermarking
Prior work on attributing LLM-generated text has largely followed two directions: post-hoc detection (Mitchell et al., 2023; Su et al., 2023; Bao et al., 2024; Zhu et al., 2023; Yang et al., 2024a) and LLM watermarking (Kirchenbauer et al., 2023; Dathathri et al., 2024; Lu et al., 2024; Chen et al., 2024; Hou et al., 2024; Guo et al., 2024; Wang et al., 2025b; Park et al., 2026).
Post-hoc detection methods typically operate without altering the source model, making them suitable for retrospective auditing. DetectGPT (Mitchell et al., 2023) exploits probability curvature in the likelihood landscape, with subsequent work improving efficiency via log-rank–based estimators (Su et al., 2023) and conditional probability curvature (Bao et al., 2024). Other approaches leverage LLM querying through revision-based similarity analysis (Zhu et al., 2023) or analyze divergent continuations and n-gram distributions (Yang et al., 2024a). Despite strong empirical performance, these methods remain sensitive to domain shift and continued advances in generation quality.
LLM watermarking approaches address these limitations by embedding attribution signals directly into the decoding process. Kirchenbauer et al. (2023) introduce a foundational logit-bias watermarking framework, which subsequent work extends toward scalable deployment (Dathathri et al., 2024) and a range of design variants, including entropy-aware (Lu et al., 2024), lexical (Chen et al., 2024), semantic (Hou et al., 2024), and contextual (Guo et al., 2024) schemes. MorphMark (Wang et al., 2025b) advances this line by adaptively modulating watermark strength using the cumulative probability mass of green-list tokens to address the effectiveness–quality trade-off. WaterMod (Park et al., 2026) refines the paradigm through probability-ranked modular partitioning. Collectively, these approaches establish strong baselines while treating attribution as automated inference, with limited emphasis on human judgment.
2.2 Linguistic Profiling of AI-Generated Text
Recent work has shifted from purely algorithmic detection toward systematic linguistic characterization of LLM-generated text. Rather than framing detection as a black-box task, this line of research examines how model outputs differ from human writing across multiple levels of linguistic structure.
Zanotto and Aroyehun (2025) find that LLM-generated text often shows deeper and longer dependency structures than human writing, while human texts exhibit more diverse semantic content and greater overall variation. They further note that newer models become more human-like on syntactic depth and length metrics, shrinking earlier gaps. Within the Korean context, Park et al. (2025) identify language-specific artifacts in spacing, morphology, and punctuation that differentiate LLM-generated Korean text from human writing. These observations highlight the importance of attending to fine-grained linguistic realizations beyond surface-level fluency.
2.3 Human Expertise in AI Detection
Recent studies increasingly emphasize the role of expertise, experience, and structured feedback in shaping human detection performance. Rather than treating human judgment as inherently unreliable, this work shows that detection ability varies substantially with exposure and analytic framing.
Russell et al. (2025) demonstrate that frequent LLM users outperform many automatic detectors when identifying synthetic English articles, while providing interpretable rationales grounded in linguistic cues. Dugan et al. (2023) demonstrate that boundary detection remains challenging and exhibits substantial inter-annotator variability; nevertheless, targeted interventions are shown to produce measurable improvements over time.
3 Intuition and the Detection Gap
3.1 Experimental Setup
We use the KatFish benchmark (Park et al., 2025) and focus on argumentative essays, which require a non-trivial balance of stance, evidence, and stylistic control. The evaluation set for each phase contains 30 essays selected through stratified sampling across elementary, middle, and high school levels. Each level contains two human-written essays and eight LLM-generated essays, with two samples each from GPT-4o, Solar-10.7B, Qwen2-72B, and Llama3.1-70B.
We recruit three advanced undergraduate annotators majoring in Korean Language and Literature. This results in a small yet linguistically trained panel characterized by a high degree of sensitivity to Korean grammar, discourse, and style, together with limited prior exposure to diverse LLM outputs. In Phase 1, annotators receive no rubric or explicit decision criteria and perform blind human-versus-AI attribution one essay at a time.
For a practical zero-shot judge baseline, we additionally evaluate GPT-5.2 Thinking, Gemini 3 Flash, and Claude Sonnet 4.5 under the same blind O/X instruction. We use separate generator and judge models to avoid self-recognition confounds and to estimate practical reviewing performance from strong independently developed LLMs.
Evaluator Participant Accuracy (%)
Human Annotator 1 50.00
Annotator 2 46.67
Annotator 3 80.00
Human Majority Vote 60.00
LLMs GPT-5.2 Thinking 96.67
Gemini 3 Flash 96.67
Claude Sonnet 4.5 90.00
LLMs Majority Vote 96.67
Table 1: Detection accuracy for human annotators and LLM detectors in Phase 1.
3.2 Human–LLM Detection Gap
Table 1 shows a substantial gap (36.67%p) between intuition-only human attribution and zero-shot LLM reviewing under identical instructions. On this benchmark, strong LLM judges achieve high accuracy, whereas linguistically trained humans are much less reliable when relying on intuition alone. We do not interpret this gap as evidence that human judgment is of limited use; rather, we take it to suggest that unguided human judgment is susceptible to a fluency trap, in which polished local well-formedness can obscure deeper attribution-relevant irregularities.
3.3 Analysis: Initial Human Perception
We summarize the heuristics that annotators use in the initial blind phase before they access gold labels or a formal rubric. We organize these heuristics around two recurring families: structural regularity and discourse plausibility. Annotators treat structural rigidity—overly template-like discourse organization with unnaturally “perfect” balance—and mechanical consistency—unusually uniform surface behavior in connective usage, punctuation, and spacing with little natural drift—as primary signals of machine generation.
They further characterize machine-generated essays as exhibiting overly literal adherence to prompts, limited semantic risk-taking, and forms of argumentation that remain normatively safe while lacking substantive depth. By contrast, they associate human writing with localized fallibility, creative deviation, and more organically grounded rhetorical progression. Importantly, these pre-calibration impressions are not yet stable diagnostic criteria; they are cue candidates that Phase 2 later re-anchors through explicit scoring rules.
Category Evaluation Dimension Points
Content (40) Thesis Clarity 10
Evidence Adequacy and Sufficiency 15
Consideration of Counter-arguments 5
Creativity 10
Organization (10) Global Paragraph-Level Organization 5
Sentence-to-Sentence Cohesion 5
Expression (50) Lexical Usage 10
Sentence Construction 10
Orthographic and Grammatical Conventions 10
Writing Conventions 10
Creativity 10
Total 100
Table 2: Structured overview of the LREAD rubric with point allocation. Detailed scoring criteria for each dimension are provided in the Appendix A.
4 LREAD: A Human-Centered Cognitive Calibration Framework
4.1 LREAD Construction from Linguistic Standards and Phase 1 Calibration
We derive LREAD from the 2025 national standards for argumentative writing proficiency issued by the National Institute of Korean Language [^2] Diagnostic Criteria for Argumentative Writing Ability (논증적 글쓰기 능력 진단 기준안), National Institute of Korean Language (국립국어원). ; however, the rubric itself is developed only after the completion of the Phase 1 blind evaluation. This sequencing is central to our study design. Once annotators finished the intuition-only task and received the gold labels, we mapped their reflections and disagreement cases onto criterion-anchored writing standards. Accordingly, LREAD is not a ready-made writing rubric; rather, it is a post-Phase-1 calibration instrument grounded in both external standards and empirically observed failure modes.
Importantly, the high-level rubric dimensions—Content, Organization, and Expression—remain stable, while some micro-diagnostics are intentionally language- and genre-sensitive. In the present Korean instantiation, these include orthography, spacing, register control, and natural Korean phrasing judgments that are difficult to capture through coarse notions of fluency alone. This design preserves comparability with established writing assessment principles while increasing sensitivity to the specific artifacts that emerged in our blind human–LLM discrimination task.
We make two detection-oriented adaptations. First, we retain creativity-related criteria because human essays often depart from probabilistically typical realization through localized novelty, idiosyncratic phrasing, or non-formulaic argumentative development. Second, we increase scoring granularity so that annotators attend to subtle lexical, orthographic, register, and cohesion-level variation that modern LLMs often smooth out. The resulting rubric is designed to surface cases in which machine-generated essays appear systematically too correct, too regular, or too register-stable relative to authentic student writing.
Table 2 summarizes the structured allocation of evaluation dimensions and point weights in LREAD. The rubric assigns 100 total points across content, organization, and expression, with the largest share placed on expression-level criteria. This weighting reflects our empirical observation from the Phase 1 blind study: contemporary LLM outputs often satisfy broad coherence and structural adequacy, while finer-grained expressive variation remains more diagnostic for human authorship. Creativity is therefore included as one informative component of the rubric rather than as its sole or primary discriminator; specifically, it contributes 20 of the 100 total points across content and expression.
By combining criterion-anchored writing standards with post-hoc calibration from real annotator errors, LREAD operationalizes human-sensitive cues across multiple linguistic strata. This enables annotators to move beyond surface-level fluency and identify subtle yet systematic traces of machine generation that frequently elude automated detection.
Phase Cognitive Mode Intervention & Activity Evaluation Focus
Phase 1 Intuitive baseline Blind attribution using unstructured linguistic intuition Quantifying susceptibility to the fluency trap and eliciting early cue candidates
Phase 2 Criterion-anchored analysis Application of the LREAD rubric with explicit scoring and justification Measuring gains in accuracy, consistency, and reasoning transparency
Phase 3 Internalized application Blind attribution on a limited held-out elementary-persona subset using internalized LREAD criteria Testing application to a residual difficulty condition and examining consensus on a small held-out set
Table 3: Summary of the three-phase calibration framework. The protocol is designed to study human attribution as a learnable, auditable procedure rather than a fixed trait.
4.2 Three-Phase Cognitive Calibration
We propose a three-phase longitudinal calibration framework designed to study whether human attribution can improve through explicit feedback and structured analytic scaffolding. Rather than treating human judgment as static, the framework models attribution as a learnable procedure whose quality can be assessed through changes in accuracy, agreement, and reasoning transparency. The overall trajectory of this framework, including the transition of cognitive modes and evaluation targets at each stage, is summarized in Table 3.
Phase 1 establishes an intuitive baseline through a zero-shot blind detection task, as described in Section 3. During this phase, annotators rely solely on unstructured linguistic intuition, without access to a formal rubric or explicit decision criteria. We use this phase to probe initial failure modes, including susceptibility to surface-fluency bias, whereby stylistic polish in LLM-generated text can obscure underlying structural artifacts. Following Phase 1, we disclose the ground-truth labels and the identities of the four source models. This disclosure serves as an initial calibration signal, enabling annotators to retrospectively align their intuitive judgments with recurring output patterns and disagreement cases.
Phase 2 introduces structured analytical detection. Building on the insights gained from Phase 1, annotators collaboratively design the LREAD rubric by integrating national writing standards with empirically observed failure cases. We then conduct a second blind evaluation using a new set of 30 essay samples, stratified evenly across elementary, middle, and high school levels. We construct these samples to be disjoint from those used in Phase 1 and deliberately withhold the label distribution to mitigate memorization effects and distributional inference. During this phase, annotators justify each classification decision by providing explicit, multi-dimensional scores grounded in the LREAD rubric. This design elicits rubric-scaffolded reasoning, prompting evaluators to align perceived linguistic irregularities with formal analytic categories rather than relying on holistic impressions.
Phase 3 evaluates internalized application on a limited held-out subset. After a second disclosure of ground-truth labels, annotators perform a final blind detection task focused exclusively on elementary-level essays. Our focus on the elementary-persona condition does not stem from the assumption that it constitutes the most difficult setting in all cases; rather, it reflects the fact that this condition remained the primary residual challenge after Phase 2 calibration and therefore serves as a focused evaluation of whether the rubric can be applied without additional external scaffolding. Our evaluation set contains 10 previously unseen essays: 2 human-written and 8 LLM-generated (2 each from GPT-4o, Solar, Qwen2, and Llama3.1), sampled using the same stratified strategy as in Phase 1.
We treat this phase as a test of whether annotators can detect persona-mismatch artifacts, that is, discrepancies between the assumed student persona and the realized linguistic or discourse properties of the text. Such artifacts include essays framed as the work of an elementary student but marked by overly mature vocabulary, adult-like institutional reasoning, or an unusually polished and textbook-like organization of discourse. This framing is consistent with prior work showing that state-of-the-art LLMs struggle to control age-appropriate lexical simplicity and readability in child-oriented generation (Valentini et al., 2023), and that expert evaluators often perceive LLM-generated child-like text as linguistically more advanced than the target age group (Hassan et al., 2025). More broadly, Wang et al. (2025a) find that LLMs have difficulty imitating implicitly defined writing styles, often defaulting to structurally over-regularized and stylistically polished outputs rather than capturing subtle, informal authorial signals.
In the final phase, annotators apply the internalized rubric in a blind binary human-versus-AI attribution task, with optional uncertainty marking. Only after completing the blind task do we elicit explicit accounts of each source model’s linguistic fingerprints, alongside defining features of authentic human writing; these post-hoc reflections are used solely for qualitative analysis. Because the held-out subset is limited in size, we interpret Phase 3 as a preliminary assessment of criterion internalization rather than a definitive demonstration of robust generalization across domains.
4.3 Experimental Procedure
We maintain strict information control by withholding label distributions and using disjoint essay sets across phases. This design limits memorization and distributional inference, helping isolate calibration effects from familiarity.
Annotators progress from intuition-only screening to rubric-guided analysis and then to internalized application on a small held-out subset. This phased design allows us to assess how explicit criteria affect individual attribution decisions and panel-level consistency.
5 Calibration Outcomes
5.1 Accuracy Gains Across Calibration Phases
Figure 1 summarizes majority-vote performance across the three phases. In Phase 1, intuition-only attribution reaches 0.60 accuracy on 30 essays and is not significantly above chance, illustrating the severity of fluency-driven error before calibration. In Phase 2, rubric-guided attribution improves to 0.90 on a disjoint 30-essay set, indicating that criterion-anchored evaluation substantially improves performance. In Phase 3, the panel classifies all 10 essays correctly on a limited held-out elementary-persona subset; however, the exact 95% confidence interval [0.692, 1.000] indicates substantial residual uncertainty. We therefore interpret Phase 3 as encouraging small-set evidence of successful internalized application rather than a population-level estimate of perfect human detection.
[…]
Figure 1: Detection accuracy across the three phases for each human annotator and for the majority vote. Exact 95% confidence intervals are shown for the majority vote (Phase 1: $N=30$; Phase 2: $N=30$; Phase 3: $N=10$). The black dashed line denotes the majority-vote performance of the LLM detectors in Phase 1.
5.2 Error Dynamics and Calibration
(a) Phase 1
(b) Phase 2
Figure 2: Confusion matrices for the human-annotator majority vote in Phase 1 (intuition-only) and Phase 2 (rubric-guided), with $N=30$ essays per phase. Each cell reports the number of essays in the corresponding true-label and predicted-label pair under binary Human-versus-AI attribution.
Figure 2 shows how calibration reshapes majority-vote errors. In Phase 1, majority voting yields no false positives (Human mislabeled as AI) but a considerable number of false negatives (AI mislabeled as Human). The confusion matrix shows that all 6 human-written essays are correctly classified as Human, whereas among AI-generated essays, 12 are misclassified as Human and 12 are correctly classified as AI. This pattern suggests that surface-level fluency systematically cues human authorship, causing annotators to mislabel many highly fluent LLM-generated essays as human-written under intuition-only screening.
In Phase 2, rubric guidance sharply reduces false negatives and reallocates errors toward a smaller number of false positives. The confusion matrix records 5 true-human essays correctly classified as Human and 1 mislabeled as AI, while it labels 22 AI essays correctly as AI and misclassifies only 2 as Human. This shift indicates that the rubric successfully directs attention away from superficial polish and toward more stable diagnostic cues, improving recall on AI while introducing only limited over-correction on human essays.
This asymmetric shift is important. If calibration merely induced a generalized AI-leaning response bias, we would expect a much larger increase in false positives. Instead, false positives remain nearly unchanged, whereas false negatives on AI essays drop sharply from 12 to 2. This pattern is more consistent with a targeted correction of the fluency-trap failure mode than with blanket over-detection. The reduction in false negatives is statistically significant under Fisher’s exact test ($p=0.003$).
Figure 3: Phase 2 rubric profile by generator. The radar chart reports the mean macro scores (Content, Organization, Expression) aggregated from the rubric dimensions and normalized by each category maximum (40/10/50).
Figure 3 complements the confusion-based error analysis by visualizing Phase 2 rubric score distributions across generators. By aggregating criteria into Content, Organization, and Expression, the plot frames LREAD as a structured, multidimensional competency profile and reveals systematic source-level variation across dimensions and models. In particular, the figure shows that human essays remain comparatively weaker on formal expression-related dimensions, whereas several LLM sources cluster at higher levels of rubric conformity, reinforcing our central claim that surface correctness can be a misleading signal of human authorship.
5.3 Inter-Annotator Agreement Analysis
Phase NN Accuracy 95% CI (exact) TN / FP / FN / TP FN rate Fleiss’ $\kappa$ Unanimous
Phase 1 30 0.600 [0.406, 0.773] 6 / 0 / 12 / 12 50.0% (12/24) 0.088-0.088 5/30
Phase 2 30 0.900 [0.735, 0.979] 5 / 1 / 2 / 22 8.3% (2/24) 0.243 18/30
Phase 3 10 1.000 [0.692, 1.000] 2 / 0 / 0 / 8 0.0% (0/8) 0.820 9/10
Table 4: Phase-wise summary of majority-vote performance and inter-annotator agreement. We report accuracy with exact 95% confidence intervals, majority-vote confusion counts (TN / FP / FN / TP), the false-negative rate on AI essays, inclusive Fleiss’ $\kappa$, and the number of unanimously labeled essays. Fleiss’ $\kappa$ uses the uncertainty-aware inclusive definition that maps any non-binary response to a third category.
Table 4 summarizes phase-wise majority-vote performance together with inter-annotator agreement. We report Fleiss’ $\kappa$ using the inclusive definition that maps any non-binary response to a third category, because our protocol allows an explicit uncertainty response. This formulation is appropriate for the present setting: uncertainty is itself an observable part of the calibration process, especially in earlier phases, and the inclusive mapping preserves these responses rather than removing them from the agreement analysis.
The agreement trajectory broadly mirrors the gains in attribution performance. Phase 1 yields negative agreement ($\kappa=-0.088$), indicating systematic inconsistency under intuition-only detection. Phase 2 increases agreement to $\kappa=0.243$, suggesting that rubric-guided criteria align annotators toward more stable and shared decision boundaries. Phase 3 reaches strong agreement ($\kappa=0.820$) on the held-out calibrated set, together with near-complete unanimous labeling (9/10), consistent with a consolidated evaluative framework.
Fleiss’ $\kappa$ alone, however, does not explain why disagreement occurs. A more informative pattern is found in directional ambiguity: in Phase 1, annotators often appeared to identify similar cue families, yet assigned them to opposing labels, resulting in systematic disagreement rather than random noise. After calibration, disagreement becomes both less frequent and less directionally unstable, which suggests that the rubric helps align not only what annotators attend to, but also how they interpret those cues in attribution decisions.
We therefore interpret the rise in agreement as evidence of criterion alignment rather than simple convergence through repetition. This interpretation is also consistent with the table’s broader phase-wise pattern: as calibration proceeds, false negatives on AI essays drop sharply, overall accuracy improves, and unanimous decisions become substantially more common. Appendix B reports pairwise agreement tables, phase-level uncertainty summaries, and inferential tests.
6 Auditable Human Calibration vs. Zero-Shot LLM Reviewing
This comparison identifies which attribution cues remain inspectable and maintainable under human calibration, relative to zero-shot LLM review. Although strong LLM judges can perform well under blind prompting, their decisions are often opaque and vulnerable to provider-side drift. In contrast, LREAD makes attribution reasoning explicit, reviewable, and auditable.
We separate generator and judge models to avoid self-recognition confounds. Generators define the essay sources, and judge models provide an external zero-shot review baseline, better reflecting deployment with heterogeneous generators.
Dimension Calibrated human cues Zero-shot LLM-reviewer cues Interpretation
Content generic institutional solutions; weak personal constraint; safe argumentation; persona inconsistency generic neutrality; socially agreeable resolutions; emotionally flat closure Overlap at a macro level; humans are more persona-sensitive
Organization over-regularized scaffold; too-complete paragraph structure; standardized conclusion pivots canonical three-part structure; balanced concession patterns; “too complete” organization Strong overlap
Expression spacing regularity; optional comma habits; register stability; natural-Korean-phrasing issues; translationese micro-patterns near-zero typos; mature/report-like tone; broad fluency cues Strongest divergence in granularity
Table 5: Comparison of cues emphasized by calibrated human annotators and zero-shot LLM reviewers.
6.1 Human–LLM Comparison Protocol
This comparison brings together two different sources of evidence. Human signatures are elicited after Phase 3, once annotators complete calibration and receive gold labels. LLM signatures, by contrast, come from zero-shot reviewing elicited only after an initial blind O/X decision. We therefore interpret this analysis as a contrast between calibration-enriched human diagnostics and broadly accessible zero-shot LLM cues, rather than as a fairness-controlled contest. Exact blind-attribution prompts, post-hoc cue-elicitation prompts, and voting conventions are reported in Appendix C.
6.2 Translationese and Operational Granularity
As summarized in Table 5, human annotators and LLM reviewers overlap in macro-structural cues while diverging in operational granularity. Both groups attend to regular discourse structure and generic resolution patterns, suggesting that some signals of machine authorship remain accessible under zero-shot inspection.
The key difference is that calibrated humans treat punctuation, spacing, register stability, and natural Korean phrasing distortions as explicit diagnostic evidence, whereas LLM reviewers often subsume these patterns under broad fluency judgments. Such cues matter in practice only when they can be taught, inspected, and revised over time. The full after-calibration signature inventory is reported in Appendix D, including model-specific form-level and content-level cues.
6.3 Shared and Divergent Cues
Humans and zero-shot LLM evaluators converged on macro-regularity as a signal of machine authorship. Both groups identified standardized logic flow, mechanically balanced three-part organization, repetition, and generic agreeable conclusions as cues that a text had been generated rather than written under topic-specific or persona-specific constraints.
The main divergence concerned granularity. After calibration, human annotators relied on Korean-specific micro-patterns, including optional comma placement, punctuation after clause-linking endings, spacing at multiword boundaries, and register diagnostics associated with Sino-Korean nominalization. By contrast, LLM detectors emphasized broader properties such as coherence, balanced framing, and the generic safety of examples and solutions. Model-specific patterns sharpened this contrast: Qwen2 was associated with Sino-Korean nominalization and calque-like phrasing, whereas GPT-4o more often exhibited English-style translationese, including modifier-heavy noun phrases and frequent pronominal reference. Prior work links such artifacts to multilingual LLM training and reduced target-language naturalness (Doshi et al., 2024; Enomoto et al., 2025; Li et al., 2025; Guo et al., 2025). We therefore treat translationese-like interference as a discriminative cue family exploited by calibrated humans, and Appendix E outlines how this qualitative cue family could be operationalized in future work.
7 Conclusion
We provide pilot evidence that rubric-based calibration can improve human attribution of LLM-generated Korean argumentative essays under tightly controlled conditions. Across our three-phase study, linguistically trained annotators move from intuition-only screening to criterion-anchored analysis, with gains in accuracy, agreement, and reasoning transparency. More broadly, LREAD contributes a way of making attribution judgments explicit and auditable by turning diffuse intuitions into an evaluative practice that can be taught, inspected, and revised. In this sense, the value of the framework lies not only in improved performance within our controlled study, but also in offering a human-centered complement to automated detectors in settings where transparency, auditability, and maintainability matter.
Limitations
Annotator Scale and Homogeneity.
This study relies on three annotators with homogeneous training in Korean linguistics, which limits claims about population-level detection accuracy and generalization to non-expert or cross-disciplinary settings. We acknowledge that with only three raters and 10 items in Phase 3, our study is statistically underpowered to detect small effects or to rule out ceiling artifacts; the 100% majority-vote accuracy in Phase 3 should therefore be interpreted as an upper-bound demonstration rather than a robust population estimate.
At the same time, this constrained design reflects a deliberate methodological choice. By recruiting annotators with matched linguistic backgrounds and minimal prior LLM exposure, we isolate the causal contribution of cognitive calibration from confounds introduced by heterogeneous expertise or divergent annotation strategies. The resulting analysis prioritizes internal validity and theoretical clarity, establishing calibration as a learnable, structured property of human judgment rather than a demographic or individual-difference effect. Future work should expand this protocol to larger and more diverse annotator pools—including non-experts and cross-linguistic populations—to quantify how calibration dynamics interact with baseline expertise, prior LLM familiarity, and task incentives.
Genre and Domain Scope.
The empirical evaluation focuses exclusively on argumentative essays at the primary and secondary school levels, a genre that foregrounds discourse organization and reasoning coherence. Although this controlled setting enables a precise assessment of how rubric-based scaffolding reshapes judgments beyond surface-level fluency, it does not establish whether the same calibration trajectory generalizes to genres with distinct stylistic conventions, including conversational dialogue, narrative fiction, and technical documentation. Extending LREAD to these domains will require principled rubric adaptation and systematic revalidation.
Post-hoc Rubric Design.
The LREAD rubric was constructed after Phase 1, once annotators had access to ground-truth labels. While this post-hoc design ensures that the rubric encodes empirically grounded failure modes, it raises potential concerns about overfitting to the specific samples used in rubric development. We mitigate this risk by using entirely disjoint essay sets across phases and withholding label distributions throughout; nevertheless, independent validation on held-out corpora remains necessary to confirm generalizability.
Model and Temporal Generalization.
Our linguistic diagnostics are derived under fixed model versions (GPT-4o, Solar-10.7B, Qwen2-72B, Llama3.1-70B) and a single generation condition. As LLM capabilities evolve, some cues that currently discriminate machine-generated text—such as particular punctuation regularities or translationese patterns—may diminish or shift. Longitudinal and cross-model studies will be essential for distinguishing stable, transferable cues from transient artifacts tied to specific model generations, and for evaluating the durability of calibration effects as the detection landscape changes.
Ablation of Calibration Components.
Our three-phase protocol bundles multiple interventions: gold-label feedback, collaborative rubric design, and explicit scoring requirements. The current study does not isolate the contribution of each component; future work should include ablation conditions (e.g., feedback-only without rubric, rubric without gold-label disclosure) to clarify which elements drive the observed accuracy and agreement gains.
Appendix A Full LREAD Rubric and Detection-Oriented Adaptations
This section documents the full LREAD rubric as used in this study. We distinguish between (i) stable, framework-level dimensions that are likely to transfer across settings (Content, Organization, Expression) and (ii) adaptable micro-diagnostics that are intentionally sensitive to language, genre, and writer-persona conditions (e.g., orthography, spacing, register, and natural Korean phrasing / translationese-sensitive judgments). This distinction is central to interpreting LREAD as a human-centered calibration methodology instantiated in Korean rather than as a fixed universal rubric. The framework-level structure is designed to remain conceptually stable, whereas the micro-diagnostics are expected to be refreshed and revalidated when the target language, genre, or generator population changes.
Relative to the source writing standard, LREAD introduces four key adaptations for attribution: (i) finer expression-level granularity, (ii) explicit natural-Korean-phrasing judgments, (iii) dual creativity criteria at content and expression levels, and (iv) score anchors tailored to one-paragraph student essays rather than longer compositions. These adaptations clarify how LREAD differs from a conventional writing-assessment rubric and why it functions as an attribution-oriented human calibration instrument rather than as a generic composition-scoring tool.
Figure 4: The full LREAD rubric with anchored scoring criteria (Korean original).
Figure 5: The full LREAD rubric with anchored scoring criteria (English translation).
We include the complete Korean LREAD rubric in Figure 4 and an English translation in Figure 5. The rubric specifies anchored scoring criteria across three dimensions: Content (40), Organization (10), and Expression (50), with discrete score levels for each sub-criterion. Percentage-based anchors follow unit-normalized counting rules: lexical-level issues are measured as a proportion of eojeols (Korean spacing units), sentence-level issues are measured as a proportion of sentences, and orthographic-norm violations are measured as a proportion of words. We apply conservative deduplication when the same error pattern recurs, counting repeated instances of an identical grammatical phenomenon (e.g., particle misuse) as a single case, and treating repeated spelling errors with the same erroneous form as a single case. When a judgment requires contextual interpretation, annotators document the decision with a representative example excerpt to support auditability and consistent application.
Appendix B Performance, Statistical, and Rubric Diagnostics
This appendix consolidates quantitative support that is too detailed for the main narrative but essential for interpreting the study responsibly. It reports individual trajectories, uncertainty-aware phase-level performance, pairwise agreement, inferential statistics, and rubric-score descriptives. Together, these materials complement the main text by showing not only that calibration improves panel-level outcomes, but also how those improvements distribute across annotators, error types, and rubric dimensions.
B.1 Individual Annotator Accuracy Trajectories
Table 6 reports phasewise accuracy for each annotator and for the majority vote. These trajectories help separate panel-level improvement from idiosyncratic individual fluctuation.
Annotator Phase 1 ($N=30$) Phase 2 ($N=30$) Phase 3 ($N=10$)
A1 50.0% (15/30) 66.7% (20/30) 90.0% (9/10)
A2 46.7% (14/30) 96.7% (29/30) 100.0% (10/10)
A3 80.0% (24/30) 86.7% (26/30) 100.0% (10/10)
Majority vote 60.0% (18/30) 90.0% (27/30) 100.0% (10/10)
Table 6: Phasewise accuracy trajectories for individual annotators and the majority vote.
B.2 Majority-Vote Performance with Uncertainty Quantification
Table 7 reports phase-level outcomes with exact confidence intervals, confusion counts, agreement estimates, and unanimity counts. This table should not be read as evidence of population-level ceiling performance; rather, it shows that the strongest phase-level gain occurs from Phase 1 to Phase 2, while Phase 3 provides encouraging but small-sample evidence of internalized application.
Phase NN Accuracy 95% CI (exact) TN / FP / FN / TP Fleiss’ $\kappa$ $\kappa$ 95% CI Unanimous
1 30 0.600 [0.406, 0.773] 6 / 0 / 12 / 12 0.088-0.088 [0.254-0.254, 0.075] 5/30
2 30 0.900 [0.735, 0.979] 5 / 1 / 2 / 22 0.243 [0.030, 0.473] 18/30
3 10 1.000 [0.692, 1.000] 2 / 0 / 0 / 8 0.820 [0.071-0.071, 1.000] 9/10
Table 7: Majority-vote performance across phases with uncertainty quantification. Accuracy intervals are exact 95% binomial confidence intervals. Confidence intervals for Fleiss’ $\kappa$ are based on item-level bootstrap resampling (B=20, [^0] B=20{,}000). Because Phase 3 contains only 10 essays, its bootstrap interval for $\kappa$ is unstable and should be interpreted together with the unanimity count.
B.3 Pairwise Agreement and Directional Ambiguity
To complement Fleiss’ $\kappa$, Table 8 reports pairwise Cohen’s $\kappa$ across phases. The most informative pattern is the transition from directional opposition to alignment, especially for the Annotator 2–Annotator 3 pair.
Annotator Pair Phase 1 Phase 2 Phase 3
A1–A2 0.14 0.03-0.03 0.75
A1–A3 0.14 0.03 0.75
A2–A3 0.53-0.53 0.73 1.00
Table 8: Pairwise Cohen’s $\kappa$ across phases. The A2–A3 trajectory is consistent with directional ambiguity in Phase 1 and criterion alignment after calibration.
B.4 Statistical Tests
[…]
Test $p$-value / effect
Phase 1 majority vote vs. chance (50%) Binomial, $p=0.181$
[…]
False-negative reduction on AI essays Fisher’s exact, $p=0.003$
Human vs. LLM-judge majority vote (Phase 1) McNemar exact, $p<0.001$
Table 9: Inferential tests corresponding to the main claims in the paper.
B.5 Rubric-Score Diagnostics
Figure 6 and Tables 1012 provide a graded view of rubric-guided judgments beyond binary attribution. Across pooled evaluations, LLM essays tend to score higher on formal expression-related criteria, whereas human essays more often receive non-zero creativity scores. This inversion helps explain why intuition-only readers can be misled by surface correctness and why rubric-guided evaluation improves discrimination.
Figure 6: Distribution of mean Phase 2 LREAD scores by source group, aggregated over three annotators for each essay. Higher scores indicate stronger formal conformity to rubric criteria, not necessarily stronger evidence of human authorship.
Evaluator Elementary Middle High
Annotator 1 63.60 $\pm$ 19.59 72.90 $\pm$ 12.89 75.40 $\pm$ 8.54
Annotator 2 60.10 $\pm$ 13.77 72.00 $\pm$ 10.92 69.40 $\pm$ 7.72
Annotator 3 63.20 $\pm$ 13.18 68.60 $\pm$ 13.83 67.90 $\pm$ 5.78
All (pooled) 62.30 $\pm$ 15.31 71.17 $\pm$ 12.31 70.90 $\pm$ 7.90
Table 10: Phase 2 total LREAD scores by grade level (mean $\pm$ SD). Scores are summarized for each annotator and for pooled evaluator–essay evaluations. Each grade level contains 10 essays, yielding 30 pooled evaluations per level across three annotators.
Evaluator Human GPT-4o Solar Qwen2 Llama3.1
Annotator 1 50.67 $\pm$ 19.61 76.50 $\pm$ 4.09 74.00 $\pm$ 4.05 82.00 $\pm$ 3.90 70.00 $\pm$ 12.41
Annotator 2 55.33 $\pm$ 15.97 73.67 $\pm$ 4.27 64.67 $\pm$ 9.35 72.50 $\pm$ 10.56 69.67 $\pm$ 8.94
Annotator 3 54.33 $\pm$ 6.28 72.17 $\pm$ 12.45 64.33 $\pm$ 12.08 73.33 $\pm$ 8.33 68.67 $\pm$ 7.63
All (pooled) 53.44 $\pm$ 14.28 74.11 $\pm$ 7.70 67.67 $\pm$ 9.73 75.94 $\pm$ 8.79 69.44 $\pm$ 9.29
Table 11: Phase 2 total LREAD scores by generator (mean $\pm$ SD). Scores are summarized for each annotator and for pooled evaluator–essay evaluations. Each generator contributes 6 essays (2 per grade level), yielding 18 pooled evaluations per generator across three annotators.
Generator Content %>>0 Expression %>>0
Human 4.44 $\pm$ 5.11 44.40 1.39 $\pm$ 2.30 27.80
GPT-4o 0.56 $\pm$ 2.36 5.60 0.56 $\pm$ 1.62 11.10
Solar 0.56 $\pm$ 2.36 5.60 0.28 $\pm$ 1.18 5.60
Qwen2 2.78 $\pm$ 4.61 27.80 1.39 $\pm$ 2.87 22.20
Llama3.1 1.67 $\pm$ 3.83 16.70 0.33 $\pm$ 1.19 11.10
Table 12: Phase 2 creativity-related rubric scores by generator (mean $\pm$ SD and %>>0). The table reports pooled evaluator–essay evaluations across three annotators. Content and Expression correspond to the creativity criteria under the LREAD rubric. %>>0 denotes the percentage of evaluations with a non-zero creativity score.
Appendix C Statistical Reporting, Voting Conventions, and Prompt Protocols
This appendix documents implementation details that are important for reproducibility: voting conventions, uncertainty handling, statistical procedures, and prompt protocols for both human and LLM-based reviewing.
C.1 Voting Conventions
We use O = AI, X = Human, and $\triangle$ = uncertain throughout the annotation protocol. Individual annotators are allowed to use $\triangle$ when they judge a case to be insufficiently resolved by the available evidence.
For individual-level accuracy, $\triangle$ is counted as incorrect. This is consistent with the task objective: the blind evaluation requires a binary Human-versus-AI decision, and uncertainty does not itself constitute a correct attribution.
For Fleiss’ $\kappa$, $\triangle$ is treated as a third category. We adopt this inclusive formulation because uncertainty is a meaningful part of the calibration process, especially in earlier phases, and excluding such responses would artificially remove difficult cases from the agreement analysis.
For majority vote, $\triangle$ counts as one vote. When all three annotators choose different categories (O / X / $\triangle$), we conservatively resolve the essay to Human (X) in order to minimize false accusation. Under this rule, uncertain responses never override a binary majority.
C.2 Statistical Procedures
Accuracy confidence intervals are reported using the exact Clopper–Pearson method. We use exact intervals because the sample sizes are small, especially in Phase 3.
Confidence intervals for Fleiss’ $\kappa$ are estimated by nonparametric bootstrap over items, with item-level resampling and B=20, [^0] B=20{,}000 bootstrap samples. We report percentile intervals. Because Phase 3 contains only 10 essays, $\kappa$ intervals in that phase are unstable and are therefore interpreted cautiously alongside unanimity counts.
[…]
For the comparison between the human majority vote and the LLM-judge majority vote on the same Phase 1 essays, we use the exact McNemar test. This comparison is appropriate because the two majority-vote systems are applied to the same underlying item set.
C.3 Human Reflection Procedure After Phase 1
After the Phase 1 blind task and gold-label disclosure, each annotator wrote short structured reflections describing (i) which cues they had relied on, (ii) which essays they had found difficult, (iii) which initial assumptions appeared unreliable after gold-label disclosure, and (iv) which recurrent artifacts seemed potentially source-specific or human-typical. These reflections were used as inputs to the collaborative Phase 2 rubric-construction process.
The purpose of this reflection step was not to optimize performance on the Phase 1 items, but to externalize previously implicit heuristics so that they could later be re-anchored to explicit scoring criteria.
C.4 Human Signature Elicitation After Phase 3
After completing the blind Phase 3 task and receiving gold labels, annotators wrote structured notes on recurring model-specific patterns and on cues associated with authentic human writing. These notes included both form-level observations (e.g., structure, punctuation, spacing, connective usage, sentence rhythm) and content-level observations (e.g., stance, evidential style, resolution patterns, persona consistency).
The resulting notes were consolidated through iterative discussion into the shared cue taxonomy reported in Appendix D. We treat these signatures as after-calibration descriptors, not as universal or immutable properties of the underlying models.
C.5 LLM-Judge Blind Attribution Prompt
For the zero-shot LLM-reviewer baseline, each model first receives a blind binary-attribution instruction. The English translation of the prompt is as follows:
“You are a reviewer who determines the author of a Korean essay. The essay is either written by a human or generated by an AI model. The writer is an elementary-, middle-, or high-school student. If the essay is AI-generated, the AI produced it at an elementary-, middle-, or high-school level. Decide whether the author is human or AI. Answer X for human and O for AI.”
This prompt is intentionally minimal. Its purpose is to estimate practical zero-shot reviewing behavior rather than optimized prompted performance.
C.6 LLM-Judge Post-hoc Cue Elicitation Prompt
After completing the blind O/X decision task, each LLM reviewer is shown the gold labels and generator identities and is then asked to summarize the characteristics it noticed for each source. The English translation of the elicitation prompt is as follows:
“Summarize the characteristics you noticed for each author type during your review. Organize the summary for Human, GPT-4o, Solar, Qwen2, and Llama3.1. Use the format <Form> and <Content>.”
We use this second-stage prompt only for qualitative comparison with the human cue inventory. It is not part of the blind detection decision itself.
C.7 Runtime Conditions for LLM Judges
LLM judges were run through provider-facing web interfaces using provider-default decoding settings. We did not manually tune temperature, top-$p$, or other generation controls. We report this setup to reflect realistic zero-shot reviewing conditions rather than a carefully optimized prompting pipeline.
The judge models are used as practical reviewing baselines, not as calibrated or task-specialized detectors. Their role in the paper is therefore contrastive: they provide a useful benchmark for what strong zero-shot reviewing can achieve, while LREAD provides an auditable framework for explicit human reasoning.
C.8 Original Korean Prompt Text
Blind attribution prompt.
너는 지금부터 한국어로 작성된 에세이의 작성자를 판별하는 검수자야. 내가 제공하는 에세이는 사람이 작성한 글이거나 AI 모델이 생성한 글이야. 이때 글을 작성한 사람은 초등학생이거나 중학생이거나 고등학생이야. AI 모델이 생성한 경우에 AI 모델은 초등학생 수준 혹은 중학생 수준 혹은 고등학생 수준으로 글을 생성했어. 내가 제공하는 글을 쓴 주체가 사람인지 AI인지 판단해줘. 사람이면 X로 대답하고 AI면 O로 대답해줘.
Post-hoc cue elicitation prompt.
너가 검수를 하면서 느낀 작성자별 특징을 정리해줘. 사람, GPT-4o, Solar, Qwen2, Llama3.1 각각 특징을 정리해줘. 아래 양식에 맞춰서 작성해줘. <글 형식> <글 내용>
Appendix D Calibrated Cue Taxonomy and Model-Specific Signatures
This section documents the after-calibration cue inventory articulated by human annotators. We treat these cues as calibrated descriptors that became operational within our controlled study, not as universal or immutable properties of specific models. The purpose of this appendix is to show how rubric-guided calibration transforms diffuse impressions into an inspectable signature inventory that can be reviewed, debated, and adapted.
D.1 Signature Elicitation Protocol
After completing the blind Phase 3 task, annotators reviewed gold labels and wrote structured reflections on (i) recurring source-specific patterns, (ii) cues associated with authentic human writing, and (iii) disagreement-triggering cases from earlier phases. These notes were then consolidated into a shared cue taxonomy through iterative discussion.
We prioritized patterns that were either independently reported by multiple annotators or that sharply differentiated sources under identical prompting and rubric conditions. We report these signatures as after-calibration descriptors: they capture cues that became operational once annotators had explicit criteria and feedback, rather than impressions formed during zero-shot inspection.
D.2 Model-Specific Signatures
Across annotators, authentic human writing was associated with localized developmental errors, uneven but grounded elaboration, and persona-consistent everyday constraints. By contrast, the model-specific patterns below summarize the recurring cues attributed to each LLM family after calibration. Table 13 should be read as an auditable inventory of learned human diagnostics, not as a permanent taxonomy of model essence.
Generator Form-level cues Content-level cues Persona / affect cues Caution / overlap note
GPT-4o standardized essay scaffold; ordinal transitions; polished sentence boundaries; optional connective-comma insertion low-risk socially agreeable claims; forward-looking prescriptions; balanced concession moves inclusive “we” voice; mature but restrained stance; weak local lived detail can overlap with carefully coached school essays; diagnose jointly with persona fidelity and punctuation micro-patterns
Solar connective-heavy sentences; narrow connective repertoire; short rhythmic clauses; repeated optional commas redundancy with thin elaboration; claim–rebuttal oscillation; performative opinion markers emotionally assertive but shallow; less stable persona control distinguish from formulaic human writing by repetition density and weak evidential progression
Qwen2 formal connective chains; nominalization-heavy diction; dense Sino-Korean lexicon; occasional spacing errors; quotation-like emphasis educational / institutional framing; collective-duty arguments; abstract justification overly mature, adult-like register for student persona; non-native or calqued phrasing strongest translationese-sensitive profile; evaluate jointly with natural-Korean-phrasing criteria
Llama3.1 highest stylistic variance; template mixing; shorter outputs; variable comma consistency; occasional grammatical awkwardness more concrete elaboration; careful counterarguments; detailed examples; lexical repetition without paraphrase comparatively vivid persona through rhetorical questions or conversational devices can appear more human-like than other generators; use multiple cues rather than any single marker
Table 13: Calibrated model-specific signatures elicited from human annotators after Phase 3. These are after-calibration descriptors, not immutable properties of the models.
D.3 Shared Detection Signatures
Beyond model-specific fingerprints, the shared cue inventory reveals several cross-model regularities that annotators consistently treated as signals of machine authorship. First, LLM outputs exhibit unusually stable surface correctness relative to student writing, including fewer orthographic errors and more controlled sentence lengths. Second, outputs tend to flatten grade-level variation: essays nominally generated at elementary, middle, and high-school levels converge toward a similar band of polish and abstraction. Third, many essays converge on generic, institution-facing remedies (e.g., education, awareness campaigns, or system improvements) rather than grounded, personally constrained resolutions.
These regularities should be interpreted as family-level hypotheses that motivate adaptation and revalidation, not as permanent signatures guaranteed to survive model updates.
Appendix E Prospective Translationese Diagnostics for Future Operationalization
Building on the calibrated cue inventory in Appendix D, we outline translationese-sensitive diagnostics that may be operationalized in future work. We do not compute these measures in the present study; instead, we include them to clarify how an impressionistic human cue can be translated into a reproducible measurement agenda.
The proposed diagnostics correspond primarily to the Expression dimension of LREAD, especially natural Korean phrasing, register stability, and connective / punctuation behavior.
E.1 Motivation and Prior Evidence
We use translationese to denote systematic source-language fingerprints and stylistic shifts that distinguish translated text from natively authored text. Prior work suggests that such patterns can be learned and propagated during model pretraining, amplified by multilingual instruction pipelines, and reflected in reduced target-language naturalness. The cue inventory in Appendix D indicates that translationese-sensitive distortions are especially salient for Qwen2 and, to a lesser degree, GPT-4o in our dataset.
E.2 Translationese Diagnostics
Grounded in the calibrated human signature inventory, we outline a set of translationese-sensitive diagnostics that can be instantiated as measurable features:
  • Nominalization Rate: the proportion of nominalized predicates and abstract nominal constructions, capturing noun-heavy discourse packaging characteristic of literal transfer.
  • Sino-Korean Lexical Density: the share of Sino-Korean lexemes and high-register vocabulary, which often co-varies with translated or formalized style.
  • Discourse Deixis and Pronominal Reference: frequencies of demonstratives and pronominal forms (e.g., this/that-type anaphora) that are associated with English-style translated discourse structuring.
  • Connective–Comma Coupling: co-occurrence patterns between sentence-initial discourse connectives and optional comma insertion, reflecting punctuation regularities that annotators flagged as model-specific.
  • Sentence-Level Naturalness Scoring: complementary global measures of perceived naturalness, following prior work on multilingual naturalness evaluation.
These diagnostics are designed to shift translationese from an impressionistic descriptor to a reproducible family of cues. The goal is not to claim that translationese has already been quantitatively validated here, but to document a plausible path from calibrated human perception to future operational measurement.